OpenStudy (anonymous):

determine whether the given values are roots of the equation or not 3x^2+5x+2=0,x=3,x=1.

OpenStudy (lgbasallote):

we can verify by substituting x = 3 3(3)^2 + 5(3) + 2 = 0 is this equal to zero? we'll see... 3(9) + 15 + 2 27 + 15 + 2 32 + 2 34 \(\ne\) 0 let's see x = 1 3(1)^2 + 5(1) + 2 = 0 is it equal to zero? 3(1) + 5 + 2 3 + 5 + 2 8 + 2 10 \(\ne\) 0

OpenStudy (anonymous):

soooo we finished

OpenStudy (anonymous):

no

OpenStudy (anonymous):

k

OpenStudy (anonymous):

well we are finished but the anwer is no I just saying

OpenStudy (anonymous):

kk